A variable X has a mean of 185.5 and a standard deviation of 21.4. A random sample of size 67 is selected. Assuming X is normally distributed, what is the probability, to two decimal places, that the sample mean exceeds 200? Give your answer to two decimal places in the form x.xx
Answer by ewatrrr(24785)

P(z > 4.5/21.4/sqrt(67) = P(z > 4.5/2.6144) P(z > 1.721) = 0.04